OSLO COMPANY PREPARED THE FOLLOWING CONTRIBUTION FORMAT INCOME STATEMENT BASED ON A SALES VOLUME OF 1,000 UNITS THE RELEVANT RANGE OF PRODUCTION OF 500 UNITS TO 1500 UNITS):

SALES: $20,000

VARIABLE EXPENSES 12,000

CONTRIBUTION MARGIN 8,000

FIXED EXPENSES 6,000

NET OPERATING INCOME 2,000

SHOW ALL WORK

1. IF THE VARIABLE COST PER UNIT INCREASES BY $1, SPENDING ON ADVERTISING INCREASES BY $1,500, AND UNIT SALES INCREASE BY 250 UNITS, WHAT WOULD BE NET OPERATING INCOME?

2. WHAT IS THE BREAK EVEN POINT IN UNIT SALES?

3. WHAT IS THE BREAK EVEN POINT IN DOLLAR SALES?

4. HOW MANY UNITS MUST BE SOLD TO ACHIEVE A TARGET PROFIT OF $5,000?

5. WHAT IS THE DEGREE OF OPERATING LEVERAGE?

6. USING THE DEGREE OF OPERATING LEVERAGE, WHAT IS THE ESTIMATED PERCENT INCREASE IN NET OPERATING INCOME OF A 5% INCREASE IN SALES?

7. ASSUME THE AMOUNT OF THE COMPANY'S TOTAL VARIABLE EXPENSES AND TOTAL FIXED EXPENSES WERE REVERSED. IN OTHER WORDS, ASSUME THAT THE TOTAL VARIABLE EXPENSES ARE $6,000 AND THE TOTAL FIXED EXPENSES ARE $12,000. UNDER THIS SCENARIO AND ASSUMING THAT TOTAL SALES REMAIN THE SAME, WHAT IS THE DEGREE OF OPERATING LEVERAGE?

8. USING THE DEGREE OF OPERATING LEVERAGE THAT YOU COMPUTED IN THE PREVIOUS QUESTION, WHAT IS THE ESTIMATED PERCENT INCREASE IN NET OPERATING INCOME OF A 5% INCREASE IN SALES?
Answer #1

Variable cost per unit and total fixed cost remains constant within the relevant range

1.Net operating income = (20-13)*1250 – 6000-1500

= $1,250

2.Break even point in unit sales = Fixed costs/Contribution margin per unit

= 6000/8

= 750 units

3.Break even point in dollar sales = 750*20 = $15,000

4.Units required to be sold = (Target profit+ Fixed cost)/contribution margin per unit

= (5000+6000)/8

= 1375 units

5.Degree of operating leverage = contribution margin/net income

= 8000/2000

= 4 times

6.Increase in income = % increase in sales*degree of operating leverage

= 5%*4

= 20% increase

7.Degree of operating leverage = 14000/2000

= 7 times

8.% increase = 7*5 = 35%